# bayesOptimalControl.bib

### v1

@@ 5.2. Numerical scheme for HJB equation
where the function $\mathcal{H}$ is called the numerical Hamiltonian. Theorem 1 in [crandall@1984] guarantees the convergence, provided that the scheme $G$ is monotone and consistent.
@@
[…]
where $\phi(w)=\frac{1}{\sqrt{2\pi}}e^{-\frac{w^{2}}{2}}$. With a change of variable $-\frac{z}{\sqrt{z+\epsilon^{2}}}w=\tau$ we can rewrite the integral as
[…]
@@
(5.16) $\hat{\phi}(\tau)=\frac{\sqrt{z+\epsilon^{2}}}{z}\frac{1}{\sqrt{2\pi}}e^{-\frac{\tau^{2}(z+\epsilon^{2})}{2z^{2}}}$
The integral (5.15) is a convolution for each given $z$ and can be computed using Fourier transform. Note that the points $(\cdot,\cdot,\frac{z\epsilon^{2}}{z+\epsilon^{2}})$ are off grid $\mathcal{G}$, therefore the value function at these points can be interpolated from the nearby points.
5.4. Control under perfect observation
@@
[…]
@@
[…]
5.5. Simulations.
@@
(5.27) $\theta=0.25,b=0.5,C=1,\epsilon=0.9$
We truncated the domain with $m\in[-1,1]$, $z\in[0,1]$, and $T\in[0,1]$, as described in section 5.1. The step sizes $\Delta t=0.0125$, $\Delta m=0.1$, $\Delta z=0.1$ were chosen to satisfy condition 5.12.
@@
Figure 2. The value function of control under no observations problem. (A) The value function at initial time; (B) the slice of value function w.r.t. $m$ and $t$, fixed $z=1$; (C) the slice of value function w.r.t. $z$ and $t$, fixed $m=1$.
@@
Figure 3 shows the same slices of the value function in the case with noisy observations made with time step $\Delta t_{obs}=20\Delta t$. Overall we see that the cost is smaller compared to the case with no observations (Figure 2). Furthermore, according to (5.1b), at each observation point $t^{-}_{n}$ and all $z$ the value function is updated from the value function at points with significantly smaller variance $\frac{z\epsilon^{2}}{z+\epsilon^{2}}$ at time $t_{n}$, where the value function is smaller. That explains the jumps at each observation point which can be seen on Figure 3 (B) and (C). The information gained due to observation improves our knowledge and facilitates better decision making.
@@
Figure 3. The value function of control under noisy observations problem. (A) The value function at initial time; (B) the slice of value function w.r.t. $m$ and $t$, fixed $z=1$; (C) the slice of value function w.r.t. $z$ and $t$, fixed $m=1$.
@@
Figure 4 shows the optimal paths of the mean $m^{*}(t)$ and variance $z^{*}(t)$. The optimal paths are computed according to the dynamics (5.28) between the observation points and updated at the observation points using the simulated noisy observations of $X(t_{n+1})$, namely $Y(t_{n+1})~{}\sim~{}\mathcal{N}(m^{*}(t^{-}_{n+1}),z^{*}(t^{-}_{n+1})+\epsilon^{2})$ as follows:
(5.28) $\begin{cases}dm^{*}(t)=\Big{(}-\theta m^{*}(t)+\frac{\partial\hat{U}}{\partial m}(t,m^{*}(t),z^{*}(t))\Big{)}dt,&t_{n}\leq t<t_{n+1},~{}n=0,...,N\\ m^{*}(t_{n+1})=m^{*}(t^{-}_{n+1})+\frac{z^{*}(t^{-}_{n+1})}{z^{*}(t^{-}_{n+1})+\epsilon^{2}}(Y(t_{n+1})-m^{*}(t^{-}_{n+1})),&n=0,...,N\\ m^{*}(t_{0})=m_{0},\\ dz^{*}(t)=\Big{(}-2\theta z^{*}(t)+b^{2}\Big{)}dt,&t_{n}\leq t<t_{n+1},~{}n=0,...,N\\ z^{*}(t_{n+1})=\frac{z^{*}(t^{-}_{n+1})\epsilon^{2}}{z^{*}(t^{-}_{n+1})+\epsilon^{2}},&n=0,...,N\\ z^{*}(t_{0})=z_{0}\end{cases}$
@@
The measurements $Y(t_{n+1})$ are independent in the sense that $Y(t_{n+1})=m^{*}(t^{-}_{n+1})+\sqrt{z^{*}(t^{-}_{n+1})+\epsilon^{2}}~{}W_{n+1}$ with $\{W_{n+1}\}_{n\geq 0}$ i.i.d. $\sim\mathcal{N}(0,1)$. Note also that $m^{*}$ is updated using the obtained observation, while $z^{*}$ evolves deterministically. The probability density function, $\rho(x)$, over the optimal path becomes more peaked around the mean from one observation point to the next as we collect information through Bayesian updates, as seen in Figure 4(C).
@@
Figure 4. The optimal paths of the mean $m^{*}(t)$ and variance $z^{*}(t)$ of control under noisy observations problem. (A) The optimal paths of the mean $m^{*}(t)$ for several simulated scenarios; (B) the optimal path of the variance $z^{*}(t)$; (C) the probability density function, $\rho(x)$, over one of the optimal paths (the blue line on plot (A)).
@@
Figure 5. The comparison of the value functions at time $t=0$ in three cases. (A) The value function of problem with full information; this case corresponds to the variance $z=0$ in cases B and C with partially available information, and $x$ is the exact known state which we compare to the mean $m$ in B and C; (B) the slice of the value function of the problem with observations w.r.t. $m$ and $t$, fixed $z=1$; (C) the slice of the value function of the problem with no observations w.r.t. $m$ and $t$, fixed $z=1$.
@@ 6. Multidimensional example - Kalman filters with control
The sets of summation indices above are =$\mathcal{I}=\{1,...,d\}$, $\mathcal{J}_{i}=\{i,...,d\}$, and we will later use the notation $\mathcal{J}^{+}_{i}=\{i+1,...,d\}$ as well. Following Remark 3.12, since we are working with multivariate Gaussian distributions, it is enough to choose $\varphi_{1,i}(x)=x_{i}$, $\varphi_{2,ij}(x)=x_{i}x_{j}$, $i,j=1,...,d$, and then represent the value function
@@
(6.10) and the Kalman gain matrix $\text{and the Kalman gain matrix }K_{t_{i}}=\Sigma_{t^{-}_{i}}H^{T}\Big{(}H\Sigma_{t^{-}_{i}}H^{T}+\epsilon^{2}I\Big{)}^{-1}$
@@
Summing up, in this section we have characterized the different ingredients to set up HJB equation in our multivariate Gaussian process case. The Hamiltonian defined by (6) determines the evolution of the value function between observation times $t_{i-1}$, $t_{i}$, $i=1,...,n$, and the conditional expectation at each of the observation times $t_{i}$ (6.11) using the Kalman filter equations. It is important to emphasize at this point that the state dimension of the time dependent HJB PDE is $d+\frac{d(d+1)}{2}$, which is relatively large, even for $d=2$. This means that one should be particularly careful when choosing appropriate discretization tools, to address the ensuing curse of dimensionality.

### v2

@@ 5.1.2. Numerical scheme for HJB equation
where the function $\mathcal{H}$ is called the numerical Hamiltonian. Theorem 1 in [CL84] guarantees the convergence, provided that the scheme $G$ is monotone and consistent.
@@
(5.14) $\widehat{U}(t_{n}^{-},m,z)=\int_{\mathbb{R}}\widehat{U}\Big{(}t_{n},m+\frac{z}{\sqrt{z+\varepsilon^{2}}}w,\frac{z\varepsilon^{2}}{z+\varepsilon^{2}}\Big{)}\phi(w)dw,$
where $\phi(w)=\frac{1}{\sqrt{2\pi}}e^{-\frac{w^{2}}{2}}$. With a change of variable $\frac{z}{\sqrt{z+\varepsilon^{2}}}w=\tau$ we can rewrite the integral as
(5.15) $\widehat{U}(t_{n}^{-},m,z)=\int_{\mathbb{R}}\widehat{U}\Big{(}t_{n},m-\tau,\frac{z\varepsilon^{2}}{z+\varepsilon^{2}}\Big{)}\hat{\phi}(\tau)d\tau,$
@@
(5.16) $\hat{\phi}(\tau)=\frac{\sqrt{z+\varepsilon^{2}}}{z}\frac{1}{\sqrt{2\pi}}e^{-\frac{\tau^{2}(z+\varepsilon^{2})}{2z^{2}}}$
The integral (5.15) is a convolution for each given $z$ and can be computed using Fourier transform. Note that the points $(\cdot,\cdot,\frac{z\varepsilon^{2}}{z+\varepsilon^{2}})$ are off grid $\mathcal{G}$, therefore the value function at these points can be interpolated from the nearby points.
5.1.4. Control under perfect observation
@@
(5.23) $H(x,t,\partial_{x}u,\partial_{x}^{2}u)$ $=\underset{\alpha}{\min}\{(-\theta x+\alpha)\partial_{x}u+b^{2}\partial^{2}_{x}u/2+x^{2}+C\alpha^{2}\}$
@@
and $g(t)=\sigma^{2}\int_{t}^{T}f(s)ds$. Equation (5.25) can be integrated numerically. Thus, computing $u(t,x)$ provides a lower bound on the value function for the case with noisy measurements.
5.1.5. Simulations.
@@
(5.27) $\theta=0.25,b=0.5,C=1,\varepsilon=0.9$
We truncated the domain with $m\in[-1,1]$, $z\in[0,1]$, and $T\in[0,1]$, as described in section 5.1.1. The step sizes $\Delta t=0.0125$, $\Delta m=0.1$, $\Delta z=0.1$ were chosen to satisfy condition 5.12.
@@
Figure 2. The value function of linear-quadratic control problem under no observations. (A) The value function at initial time; (B) the slice of value function w.r.t. $m$ and $t$, fixed $z=1$; (C) the slice of value function w.r.t. $z$ and $t$, fixed $m=1$.
@@
Figure 3 shows the same slices of the value function in the case with noisy observations made with time step $\Delta t_{obs}=20\Delta t$. Overall we see that the cost is smaller compared to the case with no observations (Figure 2). Furthermore, according to (5.1b), at each observation point $t^{-}_{n}$ and all $z$ the value function is updated from the value function at points with significantly smaller variance $\frac{z\varepsilon^{2}}{z+\varepsilon^{2}}$ at time $t_{n}$, where the value function is smaller. That explains the jumps at each observation point which can be seen on Figure 3 (B) and (C). The information gained due to observation improves our knowledge and facilitates better decision making.
@@
Figure 3. The value function of linear-quadratic control problem under noisy observations. (A) The value function at initial time; (B) the slice of value function w.r.t. $m$ and $t$, fixed $z=1$; (C) the slice of value function w.r.t. $z$ and $t$, fixed $m=1$.
@@
Figure 4 shows the optimal paths of the mean $m^{*}(t)$ and variance $z^{*}(t)$. The optimal paths are computed according to the dynamics (5.28) between the observation points and updated at the observation points using the simulated noisy observations of $X(t_{n+1})$, namely $Y(t_{n+1})~\sim~\mathcal{N}(m^{*}(t^{-}_{n+1}),z^{*}(t^{-}_{n+1})+\varepsilon^{2})$ as follows:
(5.28) $\begin{cases}dm^{*}(t)=\Big{(}-\theta m^{*}(t)+\frac{\partial\hat{U}}{\partial m}(t,m^{*}(t),z^{*}(t))\Big{)}dt,&t_{n}\leq t<t_{n+1},~n=0,...,N\\ m^{*}(t_{n+1})=m^{*}(t^{-}_{n+1})+\frac{z^{*}(t^{-}_{n+1})}{z^{*}(t^{-}_{n+1})+\varepsilon^{2}}(Y(t_{n+1})-m^{*}(t^{-}_{n+1})),&n=0,...,N\\ m^{*}(t_{0})=m_{0},\\ dz^{*}(t)=\Big{(}-2\theta z^{*}(t)+b^{2}\Big{)}dt,&t_{n}\leq t<t_{n+1},~n=0,...,N\\ z^{*}(t_{n+1})=\frac{z^{*}(t^{-}_{n+1})\varepsilon^{2}}{z^{*}(t^{-}_{n+1})+\varepsilon^{2}},&n=0,...,N\\ z^{*}(t_{0})=z_{0}\end{cases}$
@@
The measurements $Y(t_{n+1})$ are independent in the sense that $Y(t_{n+1})=m^{*}(t^{-}_{n+1})+\sqrt{z^{*}(t^{-}_{n+1})+\varepsilon^{2}}~W_{n+1}$ with $\{W_{n+1}\}_{n\geq 0}$ i.i.d. $\sim\mathcal{N}(0,1)$. Note also that $m^{*}$ is updated using the obtained observation, while $z^{*}$ evolves deterministically. The probability density function, $\rho(x)$, over the optimal path becomes more peaked around the mean from one observation point to the next as we collect information through Bayesian updates, as seen in Figure 4(C).
@@
Figure 4. The optimal paths of the mean $m^{*}(t)$ and variance $z^{*}(t)$ of linear-quadratic control under noisy observations problem. (A) The optimal paths of the mean $m^{*}(t)$ for several simulated scenarios; (B) the optimal path of the variance $z^{*}(t)$; (C) the probability density function, $\rho(x)$, over one of the optimal paths (the blue line on plot (A)).
@@
Figure 5. The comparison of the value functions at time $t=0$ in three cases. (A) The value function of of linear-quadratic problem with full information; this case corresponds to the variance $z=0$ in cases B and C with partially available information, and $x$ is the exact known state which we compare to the mean $m$ in B and C; (B) the slice of the value function of linear-quadratic problem with observations w.r.t. $m$ and $t$, fixed $z=1$; (C) the slice of the value function of linear-quadratic problem with no observations w.r.t. $m$ and $t$, fixed $z=1$.
5.2. Example 2: Optimal control of measurement noise
In this example, we extend our formulation to additionally control the noise of the observations $\boldsymbol{\beta}$.
5.2.1. Experiment set up
. Rather than fixing the measurement noise at $\varepsilon$, we allow each of $N$ observations
(5.29) $Y(t_{n})=X(t_{n})+\beta_{n}Z_{n},\quad\quad Z_{n}\sim\mathcal{N}(0,1),\quad n=1,\dots,N,$
to use an adaptive noise level $\beta_{n}\in(0,\beta_{\max}]$. Reducing $\beta_{n}$ improves accuracy but incurs an experimental cost $\frac{c_{n}}{\beta_{n}}$.
As discussed in Section 4, this creates a nontrivial trade-off between estimation error and measurement expense. The HJB associated with such value function is now given by
$-\frac{1}{4C}\left[\frac{\partial\widehat{U}}{\partial m}(t,m,z)\right]^{2}=0,\quad t_{n}\leq t<t_{n+1},\quad n=0,\dots,N$
(5.30b) $\widehat{U}(t_{n}^{-},m,z)$ $=\underset{\beta_{n}\in(0,\beta_{\max}]}{\min}\int_{\mathbb{R}}\widehat{U}\biggl{(}t_{n},m+\frac{z}{\sqrt{z+\beta_{n}^{2}}}w,\frac{z\beta_{n}^{2}}{z+\beta_{n}^{2}}\biggr{)}\phi(w)dw+\frac{c_{n}}{\beta_{n}},$
(5.30c) $\widehat{U}(T,m,z)$ $=m^{2}+z,$
where at each observation step $n$ we optimize components of $\boldsymbol{\beta}=\{\beta_{1},...,\beta_{N}\}$ to achieve the smallest cost.
To quantify the benefit of noise control, we compare against the fixed-noise case $\beta_{n}\equiv\varepsilon$. To compute the value function in a fixed-noise case, we replaces the minimization in (5.30b) by the single update
(5.31) $\widehat{U}(t_{n}^{-},m,z)=\int_{\mathbb{R}}\widehat{U}\biggl{(}t_{n},m+\frac{z}{\sqrt{z+\varepsilon^{2}}}w,\frac{z\varepsilon^{2}}{z+\varepsilon^{2}}\biggr{)}\phi(w)dw+\frac{c_{n}}{\varepsilon},$
5.2.2. Simulation and results
. We consider a simulation over $t\in[0,1]$ with $N=3$ measurements at
(5.32) $t_{\rm obs}=[0.25,0.5,0.75],\quad c_{n}=[0.05,0.01,0.001].$
In principle, the measurement noise level $\beta_{n}$ is unbounded, since it is modeled as Gaussian. However, in practice, real-world measurement systems have finite precision. To reflect this, we introduce an effective upper bound $\beta_{\max}$ based on the natural variability of the OU process. Over the short time interval $T-t_{0}$, the conditional variance of the OU process grows approximately as $z_{0}^{2}(T-t_{0})$. Applying the three-sigma rule which captures about 99.7% of Gaussian outcomes, we set
(5.33) $\beta_{\max}=3\sqrt{\mathbb{V}[X_{T}-X_{t_{0}}]}\approx 3z_{0}\sqrt{T-t_{0}},$
so that the maximum measurement noise remains on the same high-probability scale as the process noise. For the numerical experiments, we truncate the computational domain to $m\in[-7,7]$, $z\in[0,1]$, ensuring that it comfortably covers this range of high-probability fluctuations. Finally, we choose $\Delta t=0.0125$, $\Delta m=0.1$, $\Delta z=0.1$ to ensure the monotonicity condition (Condition 5.12) is satisfied.
Figure 6 shows slices of the value function for the case with controlled noise $\boldsymbol{\beta}$. In Figure 7 we compare the expected cost-to-go: the controlled noise strategy always outperforms any fixed-noise baseline $\varepsilon$. Figure 8 shows the typical optimal paths of the mean $m^{*}(t)$, the variance $z^{*}(t)$, and controls $\alpha^{*}(t)$ and $\boldsymbol{\beta}^{*}$. The average optimal noise levels are given by $\boldsymbol{\beta}^{*}=[0.49,0.52,0.75]$ (estimated over $10^{4}$ simulations). Figure 9 shows the optimal experiment noises $\beta_{1}^{*}$, $\beta_{2}^{*}$ and $\beta_{3}^{*}$ as functions of $m$ and $z$. In the high uncertainty region (large $z$) of Figure 9(a), $\beta_{1}^{*}$ is driven to its minimum — intuitively, when uncertainty is large we invest in a very precise (low-noise) measurement. As we start the simulation with large initial variance $z_{0}=1$, we observe that a precise measurement with noise $\beta_{1}^{*}=0.49$ is chosen at first observation time $t_{1}=0.25$. This reduces the variance $z^{*}(t)$ significantly, and by the time we reach the third observation at $t_{3}=0.75$, $\beta_{3}^{*}$ rises, reflecting the decreasing value of further measurement noise reduction.
Figure 6. The slices of the value function for optimal control of measurement noise $\boldsymbol{\beta}$ problem. (A) $(m,z)$-slice at time $t_{0}=0$; (B) $(t,m)$-slice at fixed $z=1$; (C) $(t,z)$-slice at fixed $m=1$.
Figure 7. Comparison of the expected cost to go for cases with controlled noise level $\boldsymbol{\beta}^{*}$ and fixed noises $\varepsilon$ (averaged over $10^{4}$ simulations).
Figure 8. (A) The optimal path of the mean $m^{*}$; (B) the optimal path of the variance $z^{*}$; (C) control $\alpha^{*}$ over the optimal path; (D) the optimal noise levels $\beta_{n}^{*}$ at each observation step for some realizations of the process.
Figure 9. The optimal solution $\boldsymbol{\beta}^{*}$ as a function of $m$ and $z$ at three observations: (A) $t=0.25$; (B) $t=0.5$; (C) $t=0.75$.
5.3. Example 3: Optimal control with high-penalty region
To conclude this section, we illustrate our method on a one-dimensional Ornstein–Uhlenbeck process subject to a high-cost “obstacle” region, described in [TK23].
5.3.1. Experiment set up.
The state $X(t)$ evolves according to a mean-reverted dynamics:
(5.34) $dX(t)=(-\theta(X(t)-p)+\alpha(t,X(t)))dt+bdW(t),~~~\forall t\in[0,T],~~X(0)=x_{0}.$
We aim to minimize the expected cost
(5.35) $J(\alpha^{*})=\underset{\alpha(\cdot)}{\min}~\mathbb{E}\Big{[}\int_{0}^{T}Q(t,X(t))+X^{2}(t)+\alpha^{2}(t,X(t))dt+10X^{2}(T)\Big{]},$
with the obstacle function $Q(t,x)$ given by
(5.36) $Q(t,x)=\begin{cases}1000,\quad\text{if }0.3\leq t\leq 0.6,~0.1\leq|x|\leq 2.0,\\ 0,\quad\quad\text{otherwise}\end{cases}$
A single noisy measurement is taken at time $t_{1}=T/2$
(5.37) $Y(t_{1})=X(t_{1})+\varepsilon Z(t_{1}),~~~Z\sim\mathcal{N}(0,1).$
It is easy to show that the HJB equation for this problem reads:
(5.38a) $\frac{\partial\widehat{U}}{\partial t}(t,m,z)+\bar{Q}(t,m,z+m^{2})+m^{2}+z-\theta(m-p)\frac{\partial\widehat{U}}{\partial m}(t,m,z)+$
(5.38b) $\quad+(b^{2}-2\theta z)\frac{\partial\widehat{U}}{\partial s}(t,m,z)-\frac{1}{4}\left(\frac{\partial\widehat{U}}{\partial m}(t,m,z)\right)^{2}=0,~~t_{n}\leq t<t_{n+1},\quad n=0,...,2$
(5.38c) $\widehat{U}(t_{n}^{-},m,z)=\int_{\mathbb{R}}\widehat{U}\biggl{(}t_{n},m+\frac{z}{\sqrt{z+\varepsilon^{2}}}w,\frac{z\varepsilon^{2}}{z+\varepsilon^{2}}\biggr{)}\phi(w)\mathrm{d}w,~~n=1$
(5.38d) $\widehat{U}(T,m,z)=10\Big{(}m^{2}+z\Big{)},$
with optimal control given in closed form:
(5.39) $\alpha^{*}(t)=-\frac{1}{2}\frac{\partial\widehat{U}}{\partial m}(t,m,z).$
5.3.2. Simulation and results.
We simulate the process over $t\in[0,1]$ with a single observation at time $t_{1}=0.5$, and the domain defined by $m\in[-8,8]$, $z\in[0,1]$, with step sizes $\Delta t=5\cdot 10^{-4}$, $\Delta m=\Delta z=0.0625$. The parameters of the OU process are $p=2$, $b=0.5$, and observation noise $\varepsilon=0.2$.
Figure 10 shows the slices of the value function. As seen in Figure 10(a), at time $t_{0}=0$ the value function has a singularity at $m=0$ and is asymmetric about the origin. The obstacle region $\{0.3\leq t\leq 0.6,~-2\leq m\leq 2\}$ appears as a clearly visible elevation in the surface of $(t,m)$-slice at $z=1$ (Figure 10(b)). Although there is a narrow gap in the obstacle for $0.1\leq x\leq 0.1$, its width is so small relative to the process variance that paths will very likely intersect the high-cost region — so the control must still push the paths around the obstacle. Figure 10(c) shows $(t,z)$-slice at $m=1$. Being at $m=1$ during $0.3\leq t\leq 0.6$ incurs high cost, which appears as an elevated band.
Figure 10. The slices of the value function for optimal control with high-penalty region. (A) $(m,z)$-slice at time $t_{0}=0$; (B) $(t,m)$-slice at fixed $z=1$; (C) $(t,z)$-slice at fixed $m=1$.
Figure 11 shows the optimal paths of the mean $m^{*}(t)$, the variance $z^{*}(t)$, optimal control $\alpha^{*}(t)$ and cost over the optimal path for initial point $(m_{0},z_{0})=(0,1)$. As seen in Figure 10(a), the optimal mean $m^{*}(t)$ is pushed far enough from the obstacle so that the probability of hitting the obstacle is low. Because of the high terminal cost $10X^{2}(T)$, the control drives the mean towards zero. After the observation, the variance $z^{*}(t)$ drops as seen from Figure 10(b), which allows the trajectories to pass closer to the obstacle with less uncertainty. Figure 10(c) shows the corresponding optimal control $\alpha^{*}(t)$. Note in the blue sample the large positive push right after $t=0.5$, needed to avoid the obstacle. The yellow path, starting further out, receives a negative push to drive the mean back towards zero.
Figure 11. The optimal paths of (A) mean $m^{*}(t)$, (B) variance $z^{*}(t)$, (C) controls $\alpha^{*}(t)$, and (D) cost to go, given the initial point $(m_{0},z_{0})=(0,1)$. The solution by HJB is $\widehat{U}(t_{0},m_{0},z_{0})=81.0159$, and estimated over optimal paths by Monte Carlo with 100 simulations is $J\approx 79.7914\pm 2.2261$ (95% CI: $[77.5653,82.0175]$).
Figure 12 shows plots of the optimal paths of the mean $m^{*}$, the variance $z^{*}$, the optimal control $\alpha^{*}$ and the cost over the optimal path for a different initial point - $(m_{0},z_{0})=(-0.125,1)$. Starting from a slightly negative initial mean $m_{0}=-0.125$, the optimal strategy pushes the mean around the obstacle on the opposite side. Otherwise, the behavior of the optimal paths are analogous to the previous case.
Figure 12. The optimal paths of (A) mean $m^{*}(t)$, (B) variance $z^{*}(t)$, (C) controls $\alpha^{*}(t)$, and (D) cost to go, given the initial point $(m_{0},z_{0})=(-0.125,1)$. The solution by HJB is $\widehat{U}(t_{0},m_{0},z_{0})=80.6807$, and estimated over optimal paths by Monte Carlo with 100 simulations is $J\approx 79.2191\pm 2.6703$ (95% CI: $[76.5488,81.8895]$).
@@ 6. Multidimensional example - Kalman filters with control
The sets of summation indices above are $\mathcal{I}=\{1,...,d\}$, $\mathcal{J}_{i}=\{i,...,d\}$, and we will later use the notation $\mathcal{J}^{+}_{i}=\{i+1,...,d\}$ as well. Following Remark 3.12, since we are working with multivariate Gaussian distributions, it is enough to choose $\varphi_{1,i}(x)=x_{i}$, $\varphi_{2,ij}(x)=x_{i}x_{j}$, i,j=1,,di,j=1,...,d, and then represent the value function
@@
(6.10) $\text{and the Kalman gain matrix }K_{t_{i}}=\Sigma_{t^{-}_{i}}H^{T}\Big{(}H\Sigma_{t^{-}_{i}}H^{T}+\epsilon^{2}I\Big{)}^{-1}$
@@
Summing up, in this section we have characterized the different ingredients to set up HJB equation in our multivariate Gaussian process case. The Hamiltonian defined by (6) determines the evolution of the value function between observation times $t_{i-1}$, $t_{i}$, i=1,,ni=1,...,n, and the conditional expectation at each of the observation times $t_{i}$ (6.11) using the Kalman filter equations. It is important to emphasize at this point that the state dimension of the time dependent HJB PDE is $d+\frac{d(d+1)}{2}$, which is relatively large, even for $d=2$. This means that one should be particularly careful when choosing appropriate discretization tools, to address the ensuing curse of dimensionality.